The mean weekly income of the trainees of an engineering firm is$1200,with a standard deviation of $250. Find the probability that an individual trainee earns more than$900 per week
Answer #1

solution:-

given that mean = 1200 , standard deviation = 250

formula z = (x - mean)/standard deviation

P(X > 900) = P(Z > (900-1200)/250)

= p(z > -1.2)

= P(Z < 1.2)

= 0.8849

note:- from z score table